How to solve this equation?

4(n+3)=2(6+2n)

Found 3 solutions by khwang, roxy9229, Cabdiqafaar:
Answer by khwang(438)

Somebody gave you a completely wrong solution.
But, that is all your own fault since it seems that you even need try
to solve this baby level question in algebra(without any single step).
Your equation is an identity just like n = n,3 =3, 5 = 5,etc
Thre are infinitely many solutions and I don't want to explain anymore.

Kenny




Answer by roxy9229(32)

4(n+3)=2(6+2n)
4n+12 = 12+4n
-4n -4n
---------------
12=12

Answer by Cabdiqafaar(9)

4(n+3)=2(6+2n)
4n+12=12+4n
4n-4n=12-12
0=0

N0TE: YOU D0N'T ASKED US THE VALUE OF (N)